### v1

@@ 1 Introduction
Quantum thermodynamics is a rapidly progressing field aimed at understanding the thermodynamics at the quantum level [1, 2]. The field is getting much attention from the scientific community not just because it provides a link between quantum mechanics and thermodynamics but also because it aids in the development of nanoscale devices that aims to harness the potential benefits of the “quantumness" in their working medium (WM) [3]. The recent advances in the experiments such as trapped ions or ultracold atoms [4, 5], using NMR techniques [6], nitrogen vacancy centres in diamond [7] have made the realization of quantum devices possible. Among the various quantum devices studied, for example, quantum refrigerators [8, 9], quantum batteries [10, 11], quantum sensors [12], quantum thermal transistors [13, 14], etc., our work focusses on quantum heat engines that follow the quantum Otto cycle [15, 16].
@@ 2 Free fermionic model
Let us consider a free fermionic model described by the Hamiltonian
@@ 5 Transverse Ising model as working medium
where $J$ is the interaction strength, $\sigma_{n}^{\mu}$ with $\mu=x,y,z$ are the Pauli matrices at site $n$, and $h$ is the transverse field which plays the role of $\alpha$ in Section 2. The model shows a quantum phase transition from the paramagnetic state ($J\ll h$) to the ferromagnetic state ($J\gg h$) at the quantum critical point $J=\pm h$ [38, 39, 40].

### v2

@@ 1 Introduction
Quantum thermodynamics is a rapidly progressing field aimed at understanding the thermodynamics at the quantum level [1, 2]. The field is getting much attention from the scientific community not just because it provides a link between quantum mechanics and thermodynamics but also because it aids in the development of nanoscale devices that aims to harness the potential benefits of the “quantumness” in their working medium (WM) [3]. The recent advances in the experiments such as trapped ions or ultracold atoms [4, 5], using NMR techniques [6], nitrogen vacancy centres in diamond [7] have made the realization of quantum devices possible. Among the various quantum devices studied, for example, quantum refrigerators [8, 9], quantum batteries [10, 11], quantum sensors [12], quantum thermal transistors [13, 14], etc., our work focusses on quantum heat engines that follow the quantum Otto cycle [15, 16, 17, 18, 19, 20].
@@ 2 Free fermionic model
For a translationally invariant system, free fermionic model can be described by the Hamiltonian
@@ 4 Universal scalings in work output
It is to be noted that Kibble Zurek scalings are valid for $L\gg\hat{\xi}\sim\tau^{\nu/\nu z+1}$ [37, 42]. Therefore, we expect the expressions given in Eqs. (22) - (25) to also hold in this limit, implying the presence of finite size corrections for small system sizes. Notably, Kibble Zurek mechanism in quantum critical systems driven out of equilibrium has been studied experimentally in quantum simulators comprising $256$ atoms [43].
@@ 5 Transverse Ising model as working medium
where $J$ is the interaction strength, $\sigma_{n}^{\mu}$ with $\mu=x,y,z$ are the Pauli matrices at site $n$, and $h$ is the transverse field which plays the role of $\alpha$ in Section 2. The model shows a quantum phase transition from the paramagnetic state ($J\ll h$) to the ferromagnetic state ($J\gg h$) at the quantum critical point $J=\pm h$ [44, 45, 46]. We set $J=1$ throughout the paper.
@@
We have assumed that the total time for a single cycle $\tau_{total}\approx\tau_{2}=\tau_{min}$. This is especially true for small $T_{C}$ when $\tau_{min}$ is large (see Eq. 25). Further the rate of evolution during a non-unitary stroke depends on the system-bath coupling strength and the system can be expected to reach infinitesimally close to thermal equilibrium in a finite time.
@@ 6 Conclusion
Acknowledgements R.B.S. and U.D. acknowledge the use of HPC facility Chandra at IIT Palakkad. U.D. acknowledges support from SERB (SPG/2022/000708). V.M. acknowledges support from SERB through MATRICS (Project No. MTR/2021/000055) and a Seed Grant from IISER Berhampur.